OpenStudy (anonymous):

Why is the following system inconsistent? t-s=1 t-s=2 2t=3s=0

OpenStudy (mrnood):

look at the first 2 equations. if a=x and a=y then x=y apply that to the first 2 equations and see what oyu get....

sammixboo (sammixboo):

Do you know what is inconsistent

sammixboo (sammixboo):

Inconsistent is a system of equations which has no solutions.

OpenStudy (skullpatrol):

The first reason is shown in the two statements: t-s=1 t-s=2 Can the difference between the same two numbers be 1 AND 2? Does 1=2? The second reason is shown in the third statement: 2t=3s=0 This statement is ONLY true for t=s=0, because we already know that ONLY 0 times any number is 0.So, comparing that to the first reason means:$$0=1=2$$

OpenStudy (ikram002p):

t-s=1 t-s=2 they are parallel :) so no solution

OpenStudy (ikram002p):



OpenStudy (anonymous):

crap that was a very stupid question... dunno how I didn't notice that... guess was the stress from the exam coming up! thank you guys !!

OpenStudy (ikram002p):

:O it was good question :) and good luck with exams !